OpenStudy (dmezzullo):

Using complete sentences, explain how to graph y = one thirdx + 4 using the slope-intercept method.

OpenStudy (dmezzullo):

DO NOT USE GOOGLE.

OpenStudy (anonymous):

start on the \(y\) axis at \((0,4)\) and then every time you go to the right 3 units, go up one unit

OpenStudy (anonymous):

You would first graph a point at (0,4) since that is the y intercept. You would then use the slope of 1/3 to graph a point on each side of the y intercept.

OpenStudy (anonymous):

We know that in slope-intercept form y = mx + b, b is a constant which represents the y-intercept. We know that y-intercepts occur when x = 0 which means that the point (0, 4) is a y-intercept therefore a point on ths graph. So start at this point, and go 1 up and 3 to the right. This is because the coefficient of x which in this case is 1/3, is always the slope in slope-intercept form. And slope = rise/run, which implies that for however many units you go up or down (rise) you then go the number of units specified by then, right or left. So here rise/run = 1/3, which means for every 1 unit up, we go 3 units to the right. And by doing that from point (0, 4), we can construct the graph. @dmezzullo